# Real-time Estimation of Bound Water Concentration during Lyophilization with Temperature-based State Observers

### v1

@@ Introduction
By using the default observer gains obtained from our simulation-based design, the observer can converge the estimated concentration to the correct value in less than 2 h for all four experiments, showing the robustness of our design (Fig. 7). In Figs. 7AB, two experiments with spatially distributed temperature measurement are considered, and so the state observer is used as usual. In Figs. 7CD, the only measurement available is the bottom temperature, and so the modified state observer is used instead (see Section 3.2). Both the original and modified observers work perfectly. The observer with spatially distributed measurement provides more complete information of the product. In addition, sensors, e.g., IR cameras, are always located outside a vial, enabling non-contact measurement. The observer with point measurement, e.g., a thermocouple, would lead to a more simple and less expensive setup, with lower noise and bias [43].
@@
Figure 7: Convergence of the estimated concentration when applying the state observer to the real systems in Case A (Panel A) and Case B (Panel B). Convergence of the estimated concentration when applying the modified state observer to the real systems in Case C (Panel C) and Case D (Panel B). In all case studies, the sampling time of 10 seconds is used, with the default parameter values and observer gains. The initial estimated concentration is chosen to by lower/higher than the true value by about 20%–100% to demonstrate convergence. The initial estimated temperature is set to the measured value.
@@
The choice of sampling time needs to be considered when implementing a state observer to the real system. If the sampling time is too large, the observer could converge slowly or diverge due to insufficient measurement information. Also, the sampling time must be higher than the computation time required for simulating the observer in each time step. Our observer can be simulated in less than a second on a normal laptop, and thus the sampling time of 10 seconds is more than adequate. The sampling time should also be chosen to be sufficiently small compared to the time scale of a process to ensure that any important dynamics are well captured. For example, the time scale of 10 seconds is reasonable given the time scale of many hours in lyophilization.
@@
The heat transfer dynamics predicted by the model is a function of the heat transfer coefficient $h$, which is estimated from data and so can have some uncertainty. For the experimental data from Cases A [15] and B [1], our observer converges the estimated states to the correct values even for 10% error in the value of the heat transfer coefficient (Fig. 8). The convergence is slightly slower, but the difference is nearly unnoticeable (cf., Figs. 7AB and 8AB). Our estimates of average concentration is insensitive to uncertainty in the heat transfer coefficient.
@@
Figure 8: Convergence of the estimated concentration when applying the state observer Case A (Panel A) and Case B (Panel B). The value of the heat transfer coefficient $h$ for the observer is set to be 10% lower than the correct value for Case A and 10% higher than the correct value for Case B. The sampling time of 10 seconds is used, with the default parameter values and observer gains. The initial estimated concentration is chosen to be different from the correct value to demonstrate convergence. The initial estimated temperature is set to the measured value.
@@ Conclusion
This article describes a new approach for the real-time estimation of bound water concentration during desorption, with application to secondary drying in lyophilization. The technique relies on a state observer, in particular a Luenberger observer, which uses the information from mechanistic understanding of the process and temperature measurement to predict the concentration of bound water. Our observer can accurately estimate the concentration for various desorption dynamics, noisy data, and real experiments. Nearly all the case studies presented in this work, except for the noise and fine-tuning parts, are achieved by a single observer design, indicating high robustness of the observer.
The proposed framework is designed to be simple and practical for implementation. The observer can be simulated in real-time, with the computation time of less than a second on a normal laptop. No concentration measurement is required; only temperature measurement is necessary. As temperature measurement is straightforward and commonly required in every step of lyophilization, the technique can be employed with a very simple setup and operation compared to any other methods. The approach is presented systematically, with detailed derivation and mathematical analysis, and so it can be easily extended to desorption-based processes other than lyophilization. This extension can be done by rewriting a mechanistic model for the new system and redesigning a state observer using the procedure described in this article.
@@
All software and data will be made available upon publication of the manuscript.

### v2

@@ D $E_{a}$ 5,920 J/mol [17]
By using the default observer gains obtained from our simulation-based design, the observer can converge the estimated concentration to the correct value in less than 2 h for all four experiments, showing the robustness of our design (Fig. 9). In Figs. 9AB, two experiments with spatially distributed temperature measurement are considered, and so the state observer is used as usual. In Figs. 9CD, the only measurement available is the bottom temperature, and so the modified state observer is used instead (see Section 3.2). Both the original and modified observers work perfectly.
The observer with spatially distributed measurement, e.g., IR cameras, provides more complete information of the product. IR cameras are always located outside a vial, enabling non-contact measurement [51, 40, 31]. However, only edge vials could be monitored, as center vials cannot be seen by the camera [52]. In this case, a model that properly incorporates the effect of thermal radiation can be used to reconstruct the temperature profiles of all vials. Noisy thermal imaging data can be filtered directly with the observer [39] or some filters such as a low-pass filter [39] and Savitzky-Golay filter [31]. The observer with point measurement, e.g., a thermocouple, would lead to a simpler and less expensive setup, with lower noise and bias [53]. This contact measurement may not be feasible for many pharmaceutical applications where contamination must be minimized. The presence of a thermocouple could also alter local heat transfer near the contact point inside the product, which is not captured by the mechanistic model; see Ref. [52] for detailed discussion on various temperature measurements in lyophilization.
@@
Figure 9: Convergence of the estimated concentration when applying the state observer to the real systems in Case A [6] (Panel A) and Case B [1] (Panel B). Convergence of the estimated concentration when applying the modified state observer to the real systems in Case C [49] (Panel C) and Case D [17] (Panel D). In all case studies, the sampling time of 10 s is used, with the default parameter values and observer gains. The initial estimated concentration is chosen to be lower/higher than the true value by about 20%–100% to demonstrate convergence. The initial estimated temperature is set to the measured value.
@@
The choice of sampling time, e.g., 10 s used in Fig. 9, needs to be considered when implementing a state observer to the real system. If the sampling time is too large, the observer could converge slowly or diverge due to insufficient measurement information. Also, the sampling time must be higher than the computation time required for simulating the observer in each time step. Our observer can be simulated in less than a second on a normal laptop, and thus the sampling time of 10 s is more than adequate. The sampling time should also be chosen to be sufficiently small compared to the time scale of a process to ensure that any important dynamics are well captured. For example, the time scale of 10 s is reasonable given the time scale of many hours in lyophilization.
@@
The heat transfer dynamics predicted by the model is a function of the heat transfer coefficient $h$, which is estimated from data and so can have some uncertainty. By using the data from Cases A [6] and B [1], we show that our observer converges the estimated states to the correct values even for 10% error in the value of the heat transfer coefficient (Fig. 10). The convergence is slightly slower, but the difference is nearly unnoticeable (cf., Figs. 9AB and 10AB). The estimated concentration is almost insensitive to uncertainty in the heat transfer coefficient in these cases.
@@
Figure 10: Convergence of the estimated concentration when applying the state observer to Case A [6] (Panel A) and Case B [1] (Panel B), with the value of the heat transfer coefficient $h$ for the observer set to be 10% lower than the correct value for Case A and 10% higher than the correct value for Case B. The sampling time of 10 s is used, with the default parameter values and observer gains. The initial estimated concentration is chosen to be different from the correct value to demonstrate convergence. The initial estimated temperature is set to the measured value.
As shown in Figs. 9 and 10, our observer can be applied for a wide range of concentration values. In secondary drying, the observer might be used for different purposes, in which the design and validation should correspond to that application. For example, the observer could be used to monitor the drying process until the moisture content is below a certain threshold. Such applications are common and do not require a highly accurate observer for the entire concentration range because only the threshold is concerned. However, there are some sensitive materials, e.g., biologics, that require precise control within a narrow range of moisture, e.g., 1%–3% [54]. In such cases, it is crucial to ensure that the observer is very accurate for the desired concentration range, meaning that the mechanistic model and temperature measurement must be highly reliable. Otherwise, using a highly accurate concentration sensor could be a better alternative.
Application to microwave lyophilization
All the results presented in the previous sections are based on conventional lyophilization, i.e., $Q_{v}=0$. In conventional lyophilization, the product contained in a glass vial is heated from the bottom via a heating shelf during the drying stages [5, 4]. To accelerate the drying process, various techniques have been studied and developed; one of the most prominent techniques is microwave lyophilization, where microwave irradiation is used to reduce the drying time during primary and secondary drying [55, 43, 44]. This section demonstrates the application of our observer for feedback control of microwave lyophilization.
Control strategies for microwave lyophilization have not been demonstrated well in the literature. It has been shown that continuous microwave heating for water desorption can lead to a huge increase in the product temperature if the microwave power is not well controlled [56]. Overheating is not desirable in lyophilization as it could lead to some serious issues such as product collapse and liquid-phase formation [4]. As such, the main goal of our feedback control system is to control the microwave power $Q_{v}$ to ensure that the product temperature does not exceed its upper limit. Here we define $Q_{v}$ by using a simple proportional control relation:
$Q_{v}(t)=K(T_{\text{up}}-T_{\text{max}}(t)),$ (50)
[…]
By using the default observer gains, the observer can accurately estimate the concentration of bound water despite the presence of microwave irradiation and feedback control (Fig. 11A). In comparison to conventional lyophilization, the drying time is shortened by about 2 h. With the proposed proportional control, the product temperature is well controlled below the maximum temperature of about 313.15 K (Fig. 11B).
Figure 11: (A) Convergence of the estimated concentration for the observer-based feedback control system where the microwave power is controlled to prevent product overheating, compared to conventional lyophilization. (B) Average product temperature under controlled microwave heating. The default observer gains $L_{T}=-1$$\times$$10^{-6}$ and $L_{c}=5$$\times$$10^{-7}$ are used. Other parameters are kept at the default values.
@@ Conclusion
A novel approach for the real-time estimation of the residual moisture during secondary drying in lyophilization is presented. The technique relies on a state observer, in particular a Luenberger observer, which uses the information from mechanistic understanding of the process and temperature measurement to predict the concentration of bound water. Our observer can accurately estimate the amount of bound water for various desorption dynamics, noisy data, and real experiments. Nearly all the case studies presented in this work, except for the noise and gain-scheduling parts, are achieved by a single observer design, indicating high robustness of the observer.
The proposed framework is designed to be simple and practical for implementation. The observer can be simulated in real time, with the computation time of less than a second on a normal laptop. No concentration measurement is required; only temperature measurement is necessary. As temperature measurement is straightforward and commonly required in every step of lyophilization, the technique can be employed with a very simple setup and operation compared to any other methods. The approach is presented systematically, with detailed derivation and mathematical analysis, and so it can be easily extended to desorption-based processes other than lyophilization. This extension can be done by rewriting a mechanistic model for the new system and redesigning a state observer using the procedure described in this article.
@@
Software and data used in this work are available at https://github.com/PrakitrSrisuma/lyo-observer-secondary.
@@ Acknowledgements
Appendices
Appendix A Numerical methods
The mechanistic model is presented in Section 2. To simulate the model, various numerical methods can be employed. Since the model equations contain the flux terms and boundary conditions, the finite volume method (FVM) is selected. The spatial domain is discretized into $m$ cells of volume $A_{c}\Delta z$ as shown in Fig. A.1, where $A_{c}$ is the cross sectional area and $\Delta z$ is given by
$\Delta z=\frac{H}{m-1}.$ (A.1)
Figure A.1: Finite volume discretization of the model equations. The variables $T_{i}$ and $c_{s,i}$ denote the temperature and bound water concentration of cell $i$.
Considering each cell as a control volume, the energy balances are
$A_{c}\frac{\Delta z}{2}\rho C_{p}\frac{dT_{1}}{dt}=q_{1}A_{c}-q_{{3}/{2}}A_{c}+A_{c}\frac{\Delta z}{2}\rho_{d}\Delta H_{s}\dfrac{dc_{s,1}}{dt}+A_{c}\frac{\Delta z}{2}Q_{v},$ (A.2)
$A_{c}\Delta z\rho C_{p}\frac{dT_{i}}{dt}=q_{i-{1}/{2}}A_{c}-q_{i+{1}/{2}}A_{c}+A_{c}\Delta z\rho_{d}\Delta H_{s}\dfrac{dc_{s,i}}{dt}+A_{c}\Delta zQ_{v},\qquad\textrm{for}\ i=2,\dots,m-1,$ (A.3)
$A_{c}\frac{\Delta z}{2}\rho C_{p}\frac{dT_{m}}{dt}=q_{m-{1}/{2}}A_{c}-q_{m}A_{c}+A_{c}\frac{\Delta z}{2}\rho_{d}\Delta H_{s}\dfrac{dc_{s,m}}{dt}+A_{c}\frac{\Delta z}{2}Q_{v}.$ (A.4)
As a result, the final discretized equations are
$\frac{dT_{1}}{dt}=\frac{2k}{\rho C_{p}}\!\left(\frac{T_{2}-T_{1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{s}}{\rho C_{p}}\!\left(\dfrac{dc_{s,1}}{dt}\right)\!+\frac{Q_{v}}{\rho C_{p}},$ (A.5)
$\frac{dT_{i}}{dt}=\frac{k}{\rho C_{p}}\!\left(\frac{T_{i+1}-2T_{i}+T_{i-1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{s}}{\rho C_{p}}\!\left(\dfrac{dc_{s,i}}{dt}\right)\!+\frac{Q_{v}}{\rho C_{p}},\qquad\textrm{for}\ i=2,\dots,m-1,$ (A.6)
$\frac{dT_{m}}{dt}=-\frac{2k}{\rho C_{p}}\!\left(\frac{T_{m}-T_{m-1}}{\Delta z^{2}}\right)\!-\frac{2h}{\rho C_{p}\Delta z}(T_{m}-T_{b})+\frac{\rho_{d}\Delta H_{s}}{\rho C_{p}}\!\left(\dfrac{dc_{s,m}}{dt}\right)\!+\frac{Q_{v}}{\rho C_{p}}.$ (A.7)
For the desorption process, the discretized equations are
$\dfrac{dc_{s,i}}{dt}=Ae^{-E_{a}/RT_{i}}c_{s,i},\qquad\textrm{for}\ i=1,2,\dots,m.$ (A.8)
Here we use the finite volume method to transform the original partial differential equations (PDEs) into a system of ordinary differential equations (ODEs). This technique is known as the method of lines [38]. The system of ODEs given by (A.5)–(A.8) can be integrated by commercial ODE solvers; MATLAB’s ode15s is used in this work.
Finally, define the state $\mathbf{x}$ and manipulated variable $\mathbf{u}$ as
$\mathbf{x}=\!\begin{bmatrix}\mathbf{T}\\ \mathbf{c_{s}}\end{bmatrix}\!,$ (A.9)
$\mathbf{u}=\!\begin{bmatrix}T_{b}\\ Q_{v}\end{bmatrix}\!,$ (A.10)
where $\mathbf{T}\in\mathbb{R}^{m}$ collects the product temperatures ($T_{1},\dots,T_{n}$) and $\mathbf{c_{s}}\in\mathbb{R}^{m}$ collects the bound water concentrations ($c_{s,1},\dots,c_{s,n}$). The heating shelf temperature $T_{b}$ and microwave power $Q_{v}$ can be manipulated to control the dynamic of secondary drying. The model equations are nonlinear in $\mathbf{x}$ and linear in $\mathbf{u}$. Consequently, (A.5)–(A.8) can be written in the vector form
$\dfrac{d\mathbf{x}}{dt}=\mathbf{F(x)}+\mathbf{Bu},$ (A.11)
where $\mathbf{F}\in\mathbb{R}^{2m}$ is the corresponding nonlinear function and $\mathbf{B}\in\mathbb{R}^{2m\times 2}$ is the matrix.
Appendix B Time constant approximation
Consider a linear system of ordinary differential equations (ODEs)
$\frac{d\mathbf{x}}{dt}=\mathbf{Mx},$ (B.1)
where $\mathbf{x}$ is the state vector and $\mathbf{M}$ is the matrix of coefficients. Note that here $\mathbf{x}$ is a general state vector, which is not the same definition as (A.9). The matrix $\mathbf{M}$ can be factorized as
$\mathbf{M}=\mathbf{U\Lambda U}^{-1},$ (B.2)
where $\mathbf{\Lambda}$ is a diagonal matrix containing the eigenvalues of $\mathbf{M}$ and $\mathbf{U}$ is the matrix of the corresponding eigenvectors.******This analysis assumes that the matrix $\mathbf{M}$ is diagonalizable, which is true for the relevant matrix in this article. A more general theoretical analysis that gives similar results is based on the Jordan form for $\mathbf{M}$, and is provided in textbooks on dynamical systems analysis or state-space control. Rearranging (B.1) with (B.2) gives
$\frac{d}{dt}\!\left(\mathbf{U}^{-1}\mathbf{x}\right)\!=\mathbf{\Lambda}\!\left(\mathbf{U}^{-1}\mathbf{x}\right).$ (B.3)
For a new set of variables defined by
$\mathbf{y}=\mathbf{U}^{-1}\mathbf{x},$ (B.4)
(B.4) can be rewritten as
$\frac{d\mathbf{y}}{dt}=\mathbf{\Lambda y}.$ (B.5)
Here the system of equations is completely decoupled. For example, consider the case where $\mathbf{x}\in\mathbb{R}^{2}$, the corresponding vectors and matrices are
$\mathbf{x}=\!\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix},$ (B.6)
$\mathbf{y}=\!\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},$ (B.7)
$\mathbf{\Lambda}=\!\begin{bmatrix}\lambda_{1}&0\\ 0&\lambda_{2}\end{bmatrix},$ (B.8)
$\mathbf{U}=\!\begin{bmatrix}u_{1,1}&u_{1,2}\\ u_{2,1}&u_{2,2}\end{bmatrix}.$ (B.9)
Consequently, (B.5) becomes
$\frac{d}{dt}\!\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}\!=\!\begin{bmatrix}\lambda_{1}&0\\ 0&\lambda_{2}\end{bmatrix}\!\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},$ (B.10)
where the analytical solution is
$\begin{bmatrix}y_{1}\\[2.84544pt] y_{2}\end{bmatrix}\!=\!\begin{bmatrix}y_{1,0}\>\mathrm{exp}(\lambda_{1}t)\\[2.84544pt] y_{2,0}\>\mathrm{exp}(\lambda_{2}t)\end{bmatrix}.$ (B.11)
Multiplying both sides of (B.11) with the matrix $\mathbf{U}$ results in
$\begin{bmatrix}x_{1}\\[2.84544pt] x_{2}\end{bmatrix}\!=\mathbf{U}\!\begin{bmatrix}y_{1,0}\>\mathrm{exp}(\lambda_{1}t)\\[2.84544pt] y_{2,0}\>\mathrm{exp}(\lambda_{2}t)\end{bmatrix}.$ (B.12)
Representing the inverse of $U$ as
$\mathbf{U}^{-1}=\!\begin{bmatrix}v_{1,1}&v_{1,2}\\ v_{2,1}&v_{2,2}\end{bmatrix},$ (B.13)
the initial condition of $\mathbf{y}$ can be written as
$\begin{bmatrix}y_{1,0}\\[2.84544pt] y_{2,0}\end{bmatrix}\!=\!\begin{bmatrix}v_{1,1}x_{1,0}+v_{1,2}x_{2,0}\\[2.84544pt] v_{2,1}x_{1,0}+v_{2,2}x_{2,0}\end{bmatrix}.$ (B.14)
Insertion of (B.14) into (B.12) gives the analytical solution
$\begin{bmatrix}x_{1}\\[2.84544pt] x_{2}\end{bmatrix}\!=\!\begin{bmatrix}u_{1,1}(v_{1,1}x_{1,0}+v_{1,2}x_{2,0})\>\mathrm{exp}(\lambda_{1}t)+u_{1,2}(v_{2,1}x_{1,0}+v_{2,2}x_{2,0})\>\mathrm{exp}(\lambda_{2}t)\\[2.84544pt] u_{2,1}(v_{1,1}x_{1,0}+v_{1,2}x_{2,0})\>\mathrm{exp}(\lambda_{1}t)+u_{2,2}(v_{2,1}x_{1,0}+v_{2,2}x_{2,0})\>\mathrm{exp}(\lambda_{2}t)\end{bmatrix}.$ (B.15)
As derived in Appendix A, we consider $\mathbf{x}\in\mathbb{R}^{2m}$. In this case, the analytical solution is
$x_{i}=\sum_{p=1}^{2m}\nu_{i,p}\>\mathrm{exp}(\lambda_{p}t),$ (B.16)
with the coefficient
$\nu_{i,p}=u_{i,p}\!\left(\sum_{q=1}^{2m}v_{p,q}x_{q,0}\right).$ (B.17)
The coefficient $\nu_{i,p}$ represents the contribution of each eigenvalue $\lambda_{p}$ to the solution. By plotting the real part of $\nu_{i,p}$, the contribution of each eigenvalue can be quantified.
For the analysis of the observer, the vector $\mathbf{x}$ is replaced with the estimation error, and the matrix $\mathbf{M}$ is replaced by the matrix $\mathbf{F^{\prime}_{ref}}+\mathbf{LC}$ derived in Section 3.3, and thus $\lambda_{1},\lambda_{2},\dots,\lambda_{2m}$ are the eigenvalues of $\mathbf{F^{\prime}_{ref}}+\mathbf{LC}$. Our state of interest is the concentration, which corresponds to $x_{m+1},x_{m+2},\dots,x_{2m}$. The average coefficient for the concentration part $i=m+1,\dots,2m$ is
$\nu_{p}=\frac{1}{m}\sum_{i=m+1}^{2m}\nu_{i,p}.$ (B.18)
By using the default parameter values and observer gains given in Table 1, it is shown that $\lambda_{m}$ and $\lambda_{m+1}$ ($m=20$) have the largest contribution because the coefficients $\nu_{m}$ and $\nu_{m+1}$ are much larger than the rest by more than three orders of magnitude (Fig. B.1). Therefore, the error dynamics are governed by these two eigenvalues. Since $\lambda_{m+1}$ is smaller than $\lambda_{m}$, it suggests the approximation
$\tau=\frac{1}{\left|\mathrm{Re}(\lambda_{m+1})\right|},$ (B.19)
as described in Section 3.3.
Figure B.1: Contribution of each eigenvalue. Each coefficient $\nu_{p}$ shows the contribution of each eigenvalue $\lambda_{p}$, with $m=20$. The default parameters and observer gains are used in this plot.